OpenStudy (anonymous):

On the last day of a Shakespeare class, my English teacher asked the students which play they liked most. 9 students liked Macbeth best, while 12 students had other favorites. What is the probability that a randomly selected Shakespeare student likes Macbeth best?

OpenStudy (anonymous):

9= Macbeth 12= Had other favorites Students Altogether= 21 \[\frac{ 9 }{ 21 } Macbeth\] \[\frac{ 12 }{ 21 } Others\]

OpenStudy (anonymous):

\[\frac{ 9 }{ 21 }-\frac{ 12 }{ 21 }= \frac{ -1 }{ 7 }\]

OpenStudy (anonymous):